Let f(n) and g(n) be asymptotically positive functions. Prove or disprove each of the following conjectures....

  1. Let f(n) and g(n) be asymptotically positive functions. Prove or disprove each of the following conjectures.

    1. f(n) = O(g(n)) implies g(n) = Ω(f(n)) .
    2. f(n) = O(g(n)) implies g(n) = O(f(n)).
    3. f(n) + g(n) = Θ(min(f(n),g(n))).
Answer #1

(a) True

Proof:

By definition as f(n)=O(g(n)), there exists constant c such that f(n) <=c*g(n).

So, g(n) >= (1/c)f(n)=c1*f(n).

Hence ,

(b) False:

Proof:

Let's prove this by counter example.

Let f(n)=n and g(n)=n2.

Clearly, f(n)=O(g(n)).

But, we can not find any constant c such that g(n)<=c*f(n).

Hence, g(n)=O(f(n)) is false.

(c) False:

Proof:

Let's prove this also by counter example.

Let f(n)=n and g(n)=n2.

Now, min (f(n)+g(n))=n.

f(n)+g(n)=n+n2>=c*n, for c=1.

Hence,

But, we can not find any constant c1 such that, f(n)+g(n)=n+n2<=c1*n.

Hence, f(n)+g(n)=O(min(f(n),g(n)) is not true.

As, by definition of for , we have to prove

Hence, is false.